# mbrs: A Library for Minimum Bayes Risk Decoding

### v1

@@ Abstract
Minimum Bayes risk (MBR) decoding is a decision rule of text generation tasks that outperforms conventional maximum a posterior (MAP) decoding using beam search by selecting high-quality outputs based on a utility function rather than those with high-probability. Typically, it finds the most suitable hypothesis from the set of hypotheses under the sampled pseudo-references. mbrs is a library of MBR decoding, which can flexibly combine various metrics, alternative expectation estimations, and algorithmic variants. It is designed with a focus on speed measurement and calling count of code blocks, transparency, reproducibility, and extensibility, which are essential for researchers and developers. We published our mbrs as an MIT-licensed open-source project, and the code is available on GitHub [^1] GitHub: https://github.com/naist-nlp/mbrs,222 YouTube: https://youtu.be/4qeHpg4PTn0 , [^3] Demo: https://mbrs-demo.streamlit.app/
@@ 1 Introduction
Figure 1: Workflow overview of mbrs.
@@
Experiments on the WMT’22 En–De general translation task demonstrated that our mbrs can not only systematically compare various metrics and algorithms but also present the runtime statistics and supplementary information of output texts.
@@ 2.1 Minimum Bayes risk (MBR) decoding
where $\Pr(\bm{y}|\bm{x})$ is the true probability of $\bm{y}$ being generated from $\bm{x}$. $u\colon\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is the utility function that measures the quality or preference of generated text under the given reference text, which is defined as $\bm{h}\succeq\bm{h}^{\prime}\iff u(\bm{h},\bm{r})\geq u(\bm{h}^{\prime},\bm{r})$ where $\succeq$ denotes the preference relation. Since the true probability $\Pr(\bm{y}|\bm{x})$ is unknown, it is replaced with the output probability calculated by a text generation model $\theta$:
@@
Because enumerating all possible output texts is infeasible, the expected scores calculated by the utility function (expected utility; EU) are estimated using pseudo-references that are sampled according to the output probabilities. Let $\hat{\mathcal{R}}$ be a bag (a.k.a multiset) of sampled pseudo-references and $\mathcal{R}$ is the support set of $\hat{\mathcal{R}}$, i.e., a set of distinct elements of $\hat{\mathcal{R}}$, they are formulated as follows:
@@
or using the model-based method (Jinnai et al., 2024) based on the output probability:
@@ 2.3 Problems of existing implementations
Currently, MBR decoding has drawn attention from research communities, and although various methods have been proposed, there is no common library that includes many of the latest studies, making it hard to compare the quality and speed systematically. mbr-nmt555mbr-nmt: https://github.com/roxot/mbr-nmt is the original implementation of Monte Carlo estimation (Eikema and Aziz, 2022), but it does not support other later methods or metrics. COMET is known as an evaluation metric, but its framework includes a command for MBR decoding, comet-mbr (Rei et al., 2020; Fernandes et al., 2022). It only supports COMET as an evaluation metric and the Monte Carlo estimation as a decoding method. Vamvas and Sennrich (2024) released mbr666mbr: https://github.com/ZurichNLP/mbr which is highly integrated into huggingface’s transformers (Wolf et al., 2020), but it only comprises their work and the vanilla MBR decoding, and it cannot be combined with other frameworks like fairseq (Ott et al., 2019) and black-box large language model services.
@@ 3.1 Main components
  • Translation Edit Rate (TER) (Snover et al., 2006)
@@
MBR: Vanilla (Eikema and Aziz, 2020, 2022) all $\mathcal{O}(|\mathcal{H}||\hat{\mathcal{R}}|)$
PruneMBR: Confidence-based pruning (Cheng and Vlachos, 2023) all N/A
@@
PMBR: Low-rank matrix completion (Trabelsi et al., 2024) all $\mathcal{O}\left(\frac{|\mathcal{H}||\hat{\mathcal{R}}|}{r}\right)$
@@
In MBR decoding, we support not only Monte Carlo estimation (Eikema and Aziz, 2022) but also model-based estimation (Jinnai et al., 2024) by passing the log-probabilities of pseudo-references. These estimators can be combined with a variety of MBR decoding algorithms. Table 1 lists the implemented variants of MBR decoding. Note that the aggregation methods in RAMBR depend on each metric. mbrs implements the aggregation methods for the BLEU (DeNero et al., 2009), chrF (Vamvas and Sennrich, 2024), and COMET (Vamvas and Sennrich, 2024; Deguchi et al., 2024) that has been proposed so far [^7] Other metrics, e.g., TER, XCOMET, and BLEURT, cannot aggregate the references due to their calculation (DeNero et al., 2009; Vamvas and Sennrich, 2024; Deguchi et al., 2024).. The listed decoders can be combined with either the Monte Carlo or model-based estimations.
@@ 4 Experiments
Decoding Utility Est. bleu chrf comet bleurt kiwi
MAPϵ 23.8 50.1 75.6 54.9 74.2
@@
QE kiwi 22.5 51.8 82.1 58.4 83.6
MBR bleu MC 25.6 52.9 75.2 54.8 73.7
MB 24.2 50.5 75.5 54.8 74.1
chrf MC 24.4 54.2 76.3 56.4 75.1
MB 24.4 51.2 76.0 55.7 74.8
comet MC 24.1 52.8 83.9 58.7 79.5
MB 24.1 51.1 81.0 57.1 77.5
bleurt MC OOT OOT OOT OOT OOT
MB OOT OOT OOT OOT OOT
RAMBR bleu MC 25.6 52.9 75.2 54.9 73.8
MB 24.2 50.5 75.5 54.8 74.1
chrf MC 24.3 54.2 76.1 56.3 75.0
MB 24.4 51.1 75.8 55.6 74.7
comet MC 23.9 52.9 83.0 58.1 79.1
MB 24.3 51.0 80.1 56.6 77.0
CBMBR comet MC 23.2 52.4 83.8 58.3 79.2
MB 22.6 51.6 82.2 57.4 77.8
Prune- bleu MC 25.7 52.9 75.0 54.6 73.6
MBR MB 24.3 52.0 75.5 54.9 74.2
chrf MC 24.5 54.3 76.2 56.3 75.1
MB 24.6 52.8 76.1 55.9 75.0
comet MC 24.0 52.8 83.9 58.7 79.5
MB 24.4 52.5 82.1 57.9 78.4
bleurt MC 22.7 51.7 77.6 63.4 75.6
MB 24.1 52.2 77.3 59.4 75.9
PMBR bleu MC 25.0 52.2 74.2 54.4 72.8
MB 24.9 52.0 73.7 54.2 72.3
chrf MC 23.0 52.9 74.3 55.0 73.1
MB 22.9 52.7 74.1 54.8 72.8
comet MC 23.9 52.7 83.8 58.7 79.5
MB 23.6 51.8 82.6 57.7 78.6
bleurt MC 22.7 51.8 77.7 63.3 75.7
MB 23.1 51.5 77.2 60.8 75.8
Oracle bleu 44.4 62.8 76.1 59.5 71.6
chrf 39.5 66.6 77.7 62.0 72.9
comet 30.2 58.0 87.0 64.1 80.4
bleurt 31.0 58.2 79.7 71.0 75.0
Table 2: Experimental results on the WMT’22 En–De general MT task.
@@ 4.1 Setup
Using our mbrs, we conducted translation experiments on WMT’22 general translation task (Kocmi et al., 2022) in En–De, and evaluated the translation quality and speed. We compared various MBR decoding methods: vanilla (Eikema and Aziz, 2020, 2022), reference aggregation (DeNero et al., 2009; Vamvas and Sennrich, 2024), centroid-based aggregation (Deguchi et al., 2024), confidence-based pruning (Cheng and Vlachos, 2023), and probabilistic MBR (Trabelsi et al., 2024). In addition to MBR decoding, we compared other decoding methods, MAP decoding (MAP) and N-best reranking using a quality estimation model (QE). We also measured the upper bound of the translation quality (Oracle) by selecting translations that maximize the evaluation metric using the reference translations.
@@
We evaluated the translation quality on bleu (Papineni et al., 2002), chrf (Popović, 2015), comet (Rei et al., 2022a), bleurt (Sellam et al., 2020), and COMETkiwi (kiwi) (Rei et al., 2022b). We employed kiwi for QE reranking, and other metrics for the utility functions of MBR decoding and the oracle. The detailed setup is described in Appendix D.
@@ 4.2 Results
Table 3: The execution time when bleu is employed as the utility function.
@@
Table 4: The execution time when comet is employed as the utility function.
@@
Table 3 and 4 show the execution time in milliseconds (msec) per sentence when we used bleu and comet as the utility functions, respectively. “QE” in Table 4 shows the reference time of QE reranking with kiwi. In the tables, “Step” indicates each step of decoding, and “E2E” reports the end-to-end total time including miscellaneous processes. Both tables demonstrate that alternative algorithms of the vanilla MBR decoding improved the speed. The additional results with other utility functions are shown in Appendix E.
@@
As described in Section 3.4, mbrs can present the additional information of the output texts. We visualized the distribution of expected utility scores in the hypothesis set by using them. Figure 2 compares the distribution of the expected utility scores between the test cases that have the maximum and minimum variance in the test set. We observed that there are cases in which the expected utility varies significantly depending on each hypothesis and cases in which it does not vary much. The latter result suggests that the N-best outputs of MBR decoding are sometimes similar to each other.
@@
mbrs also returns which indices in the hypotheses are selected by MBR decoding. Figure 3 shows the empirical cumulative distribution of ranks selected by MBR decoding from descending rankings of output probabilities. In the figure, the x-axis indicates the 0-indexed rank of output probabilities in the hypothesis set, i.e., 0 is the highest probability in the hypotheses set, and the y-axis indicates the ratio of the number of selected by MBR decoding when using COMET as the utility function.
@@ 5 Conclusion
This paper describes mbrs, a library for MBR decoding, which implements various metrics and decoding methods. It is designed to ensure transparency, reproducibility, and extensibility for researchers and developers. We will continue to implement the latest metrics and decoding methods. We hope our mbrs will be used as a tool to further improve the quality of text generation models.
@@
Appendix A Limitations
Appendix B License
@@
Appendix C Links to Our Project
Appendix D Detailed Experimental Setup
@@
We measured the speeds on an NVIDIA RTX 6000 Ada GPU with a batch size of 256 and the half-precision computation for comet, kiwi, and bleurt metrics, and on the 32 core CPUs [^10] Intel (R) Xeon (R) Platinum 8160 for bleu and chrf metrics.
@@
Appendix E Additional Experiments
E.1 Execution time on the other metrics
@@
Table 6 and 7 show the execution times of MBR decoding on the WMT’22 En–De translation task when chrf and bleurt are used as the utility functions, respectively.
E.2 Additional translation experiments on other language directions
@@
De–En En–Zh Zh–En En–Ja Ja–En
Decoding Utility Est. bleu chrf comet bleurt kiwi bleu chrf comet bleurt kiwi bleu chrf comet bleurt kiwi bleu chrf comet bleurt kiwi bleu chrf comet bleurt kiwi
MAPϵ 26.0 50.8 78.4 62.0 75.4 25.1 25.0 76.4 53.0 73.3 12.1 35.0 67.2 49.5 65.4 15.1 22.6 79.2 46.1 77.4 9.0 29.2 68.0 46.5 68.5

### v2

@@ Abstract
Minimum Bayes risk (MBR) decoding is a decision rule of text generation tasks that outperforms conventional maximum a posterior (MAP) decoding using beam search by selecting high-quality outputs based on a utility function rather than those with high-probability. Typically, it finds the most suitable hypothesis from the set of hypotheses under the sampled pseudo-references. mbrs is a library of MBR decoding, which can flexibly combine various metrics, alternative expectation estimations, and algorithmic variants. It is designed with a focus on speed measurement and calling count of code blocks, transparency, reproducibility, and extensibility, which are essential for researchers and developers. We published our mbrs as an MIT-licensed open-source project, and the code is available on GitHub [^1] GitHub: https://github.com/naist-nlp/mbrs, [^2] Read the Docs: https://mbrs.readthedocs.io/en/latest/index.html,333 YouTube: https://youtu.be/4qeHpg4PTn0 , [^4] HP: https://naist-nlp.github.io/mbrs-web.
@@ 1 Introduction
Figure 1: Workflow overview of mbrs. Decoder class gets a candidate list and outputs high-quality hypotheses. It internally calls the expected_scores() method implemented in Metric classes that calculate the expected scores for each hypothesis. The expected scores are estimated with Monte Carlo (MC) or the model-based (MB) method. In the figure, the colored boxes without border lines denote the functions or methods and the square boxes with border lines denote the abstract classes. Methods that have the same color as a class indicate that they belong to the class.
@@
[…]
@@ 2.1 Minimum Bayes risk (MBR) decoding
where $\Pr(\bm{y}|\bm{x})$ is the true probability of $\bm{y}$ being generated from $\bm{x}$. $u\colon\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is the utility function that measures the quality or preference of generated text under the given reference text, which is defined as $\bm{h}\succeq\bm{h}^{\prime}\iff u(\bm{h},\bm{r})\geq u(\bm{h}^{\prime},\bm{r})$ where $\succeq$ denotes the preference relation.555 An evaluation metric that measures the quality of output texts, e.g., Bleu (Papineni et al., 2002) or Comet (Rei et al., 2020, 2022a) in the machine translation task, is often employed for the utility function $u$. Since the true probability $\Pr(\bm{y}|\bm{x})$ is unknown, it is replaced with the output probability calculated by a text generation model $\theta$:
@@
Because enumerating all possible output texts is infeasible, the expected scores calculated by the utility function (expected utility; EU) are estimated using pseudo-references that are sampled according to the output probabilities.666 We call this calculation of the approximated expected score using pseudo-references “expectation estimation”. Let $\hat{\mathcal{R}}$ be a bag (a.k.a multiset) of sampled pseudo-references and $\mathcal{R}$ is the support set of $\hat{\mathcal{R}}$, i.e., a set of distinct elements of $\hat{\mathcal{R}}$, they are formulated as follows:
@@
or using the model-based method (Jinnai et al., 2024b) based on the output probability:
@@ 2.3 Problems of existing implementations
Currently, MBR decoding has drawn attention from research communities, and although various methods have been proposed, there is no common library that includes many of the latest studies, making it hard to compare the quality and speed systematically. mbr-nmt888mbr-nmt: https://github.com/roxot/mbr-nmt is the original implementation of Monte Carlo estimation (Eikema and Aziz, 2022), but it does not support other later methods or metrics. Comet (Rei et al., 2020; Fernandes et al., 2022) is known as an evaluation metric, but its framework includes a command for MBR decoding, comet-mbr. It only supports Comet as the utility function and the Monte Carlo estimation as a decoding method. Vamvas and Sennrich (2024) released mbr999mbr: https://github.com/ZurichNLP/mbr which is highly integrated into huggingface’s transformers (Wolf et al., 2020), but it only comprises their work and the vanilla MBR decoding, and it cannot be combined with other frameworks like fairseq (Ott et al., 2019) and black-box large language model services.
@@ 3.1 Main components
  • Translation Edit Rate (Ter(Snover et al., 2006)
@@
MBR: Vanilla (Eikema and Aziz, 2020, 2022) any $\mathcal{O}(|\mathcal{H}||\hat{\mathcal{R}}|)$
PruneMBR: Confidence-based pruning (Cheng and Vlachos, 2023) any N/A
@@
PMBR: Low-rank matrix completion (Trabelsi et al., 2024) any $\mathcal{O}\left(\frac{|\mathcal{H}||\hat{\mathcal{R}}|}{r}\right)$
@@
In MBR decoding, we support not only Monte Carlo estimation (Eikema and Aziz, 2022) but also model-based estimation (Jinnai et al., 2024b) by passing the log-probabilities of pseudo-references. These estimators can be combined with a variety of MBR decoding algorithms. Table 1 lists the implemented variants of MBR decoding. Note that the aggregation methods in RAMBR depend on each metric. mbrs implements the aggregation methods for the Bleu (DeNero et al., 2009), chrF (Vamvas and Sennrich, 2024), and Comet (Vamvas and Sennrich, 2024; Deguchi et al., 2024) that has been proposed so far [^10] Other metrics, e.g., Ter, xComet, and Bleurt, cannot aggregate the references due to their calculation (DeNero et al., 2009; Vamvas and Sennrich, 2024; Deguchi et al., 2024).. The listed decoders can be combined with either the Monte Carlo or model-based estimations.
@@ 4 Experiments
En–De De–En
Decoding Utility Est. Bleu chrF Comet Bleurt Kiwi Bleu chrF Comet Bleurt Kiwi
MAPϵ 23.8 50.1 75.6 54.9 74.2 26.0 50.8 78.4 62.0 75.4
@@
QE Kiwi 22.5 51.8 82.1 58.4 83.6 23.8 50.7 82.2 63.3 81.9
MBR Bleu MC 25.6 52.9 75.2 54.8 73.7 27.2 52.1 78.5 62.0 75.3
MB 24.2 50.5 75.5 54.8 74.1 26.4 51.1 78.5 61.9 75.3
chrF MC 24.4 54.2 76.3 56.4 75.1 27.5 53.4 79.0 62.8 76.0
MB 24.4 51.2 76.0 55.7 74.8 26.8 51.9 78.8 62.5 76.0
Comet MC 24.1 52.8 83.9 58.7 79.5 25.7 51.9 83.0 63.7 78.5
MB 24.1 51.1 81.0 57.1 77.5 26.2 51.4 81.6 63.2 77.6
Bleurt MC OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
MB OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
RAMBR Bleu MC 25.6 52.9 75.2 54.9 73.8 27.3 52.1 78.6 62.1 75.4
MB 24.2 50.5 75.5 54.8 74.1 26.4 51.1 78.5 62.0 75.4
chrF MC 24.3 54.2 76.1 56.3 75.0 27.4 53.4 79.0 62.9 76.0
MB 24.4 51.1 75.8 55.6 74.7 26.9 51.9 78.9 62.6 76.0
Comet MC 23.9 52.9 83.0 58.1 79.1 26.0 51.9 82.3 63.5 78.0
MB 24.3 51.0 80.1 56.6 77.0 26.2 51.4 81.0 63.1 77.2
CBMBR Comet MC 23.2 52.4 83.8 58.3 79.2 25.1 51.4 83.0 63.3 78.3
MB 22.6 51.6 82.2 57.4 77.8 24.2 50.4 81.9 62.6 77.4
Pruning Bleu MC 25.7 52.9 75.0 54.6 73.6 27.1 52.0 78.5 61.9 75.3
MB 24.3 52.0 75.5 54.9 74.2 26.2 51.4 78.6 61.9 75.4
chrF MC 24.5 54.3 76.2 56.3 75.1 27.4 53.5 78.9 62.8 75.9
MB 24.6 52.8 76.1 55.9 75.0 26.8 52.3 78.8 62.4 75.9
Comet MC 24.0 52.8 83.9 58.7 79.5 25.8 51.9 83.0 63.7 78.4
MB 24.4 52.5 82.1 57.9 78.4 26.3 51.9 81.6 63.3 77.6
Bleurt MC 22.7 51.7 77.6 63.4 75.6 24.9 51.3 79.4 65.1 76.4
MB 24.1 52.2 77.3 59.4 75.9 26.0 51.5 79.1 63.5 76.1
PMBR Bleu MC 25.0 52.2 74.2 54.4 72.8 26.4 51.4 78.2 61.8 75.0
MB 24.9 52.0 73.7 54.2 72.3 26.4 51.5 78.0 61.6 74.7
chrF MC 23.0 52.9 74.3 55.0 73.1 26.0 52.2 78.3 62.1 75.1
MB 22.9 52.7 74.1 54.8 72.8 25.7 52.1 78.1 61.9 75.0
Comet MC 23.9 52.7 83.8 58.7 79.5 25.9 51.9 82.8 63.6 78.3
MB 23.6 51.8 82.6 57.7 78.6 25.7 51.5 82.1 63.5 77.9
Bleurt MC 22.7 51.8 77.7 63.3 75.7 25.0 51.4 79.4 65.1 76.4
MB 23.1 51.5 77.2 60.8 75.8 25.5 51.4 79.2 64.1 76.3
Oracle Bleu 44.4 62.8 76.1 59.5 71.6 46.4 63.4 80.9 66.6 74.3
chrF 39.5 66.6 77.7 62.0 72.9 43.5 66.0 81.3 67.4 74.8
Comet 30.2 58.0 87.0 64.1 80.4 34.4 58.3 86.4 68.5 78.9
Bleurt 31.0 58.2 79.7 71.0 75.0 35.7 59.0 82.5 72.6 76.3
Table 2: Experimental results on the WMT’22 general MT task in En–De and De–En.
@@ 4.1 Setup
Using our mbrs, we conducted translation experiments on the WMT’22 general translation task (Kocmi et al., 2022) in En–De and De–En, and evaluated the translation quality and speed. We compared various MBR decoding methods: vanilla (Eikema and Aziz, 2020, 2022), reference aggregation (DeNero et al., 2009; Vamvas and Sennrich, 2024), centroid-based aggregation (Deguchi et al., 2024), confidence-based pruning (Cheng and Vlachos, 2023), and probabilistic MBR (Trabelsi et al., 2024). In addition to MBR decoding, we compared other decoding methods, MAP decoding (MAP) and $N$-best reranking using a quality estimation model (QE). We also measured the upper bound of the translation quality (Oracle) by selecting translations that maximize the evaluation metric using the reference translations.
@@
We evaluated the translation quality on Bleu (Papineni et al., 2002), chrF (Popović, 2015), Comet (Rei et al., 2022a), Bleurt (Sellam et al., 2020), and CometKiwi (Kiwi(Rei et al., 2022b). We employed Kiwi for QE reranking, and other metrics for the utility functions of MBR decoding and the oracle. The detailed setup is described in Appendix C.
@@ 4.2 Results
Table 3: The execution time in the WMT’22 En–De translation task when Bleu is employed as the utility function.
@@
Table 4: The execution time in the WMT’22 En–De translation task when Comet is employed as the utility function.
@@
Table 3 and 4 show the execution time in milliseconds (msec) per sentence when we used Bleu and Comet as the utility functions, respectively, in the WMT’22 En–De translation. “QE” in Table 4 shows the reference time of QE reranking with Kiwi. In the tables, “Step” indicates each step of decoding, and “E2E” reports the end-to-end total time including miscellaneous processes. Both tables demonstrate that alternative algorithms of the vanilla MBR decoding improved the speed. The additional results with other utility functions are shown in Appendix D.
@@
As described in Section 3.4, mbrs can present the additional information of the output texts. We visualized the distribution of expected utility scores in the hypothesis set by using them. Figure 2 compares the distribution of the expected utility scores between the test cases that have the maximum and minimum variance in the test set of the WMT’22 En–De translation task. We observed that there are cases in which the expected utility varies significantly depending on each hypothesis and cases in which it does not vary much. The latter result suggests that the $N$-best outputs of MBR decoding are sometimes similar to each other.
@@
mbrs also returns which indices in the hypotheses are selected by MBR decoding. Figure 3 shows the empirical cumulative distribution of ranks selected by MBR decoding from descending rankings of output probabilities in the WMT’22 En–De translation task. In the figure, the x-axis indicates the 0-indexed rank of output probabilities in the hypothesis set, i.e., 0 is the highest probability in the hypotheses set, and the y-axis indicates the ratio of the number of selected by MBR decoding when using Comet as the utility function.
@@ 5 Conclusion
This paper describes mbrs, a library for MBR decoding, which implements various metrics and decoding methods. It is designed to ensure transparency, reproducibility, and extensibility for researchers and developers. We will continue to implement the latest metrics and decoding methods131313 This paper describes mbrs up to the version v0.1.2. The latest version, v0.1.3, now implements diverse MBR (Jinnai et al., 2024a). . Furthermore, we will support evaluation metrics other than those used in the machine translation task. We hope our mbrs will be used as a tool to further improve the quality of text generation models.
Limitations
@@
Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP21H05054 and JP23H03458.
Appendix A License
@@
Appendix B Links to Our Project
The following are links to our project pages:
Appendix C Detailed Experimental Setup
@@
We measured the speeds on an ${}^{\text{\textregistered}}$ ${}^{\text{\texttrademark}}$ 6000 Ada GPU with a batch size of 256 and the half-precision computation for Comet, Kiwi, and Bleurt metrics, and on the 32 core ${}^{\text{\textregistered}}$ ${}^{\text{\textregistered}}$ Platinum 8160 CPUs for Bleu and chrF metrics.
@@
Appendix D Additional Experiments
D.1 Execution time on the other metrics
Table 6 and 7 show the execution times of MBR decoding in the WMT’22 En–De translation task when chrF and Bleurt are used as the utility functions, respectively.
D.2 Additional translation experiments on other language directions
[…]
@@ D.2 Additional translation experiments on other language directions
En–Zh Zh–En
Decoding Utility Est. Bleu chrF Comet Bleurt Kiwi Bleu chrF Comet Bleurt Kiwi
MAPϵ 25.1 25.0 76.4 53.0 73.3 12.1 35.0 67.2 49.5 65.4
@@
QE Kiwi 25.8 25.1 83.0 55.7 82.8 15.4 46.6 76.9 55.0 77.7
MBR Bleu MC 29.8 27.8 77.3 53.8 74.2 17.2 46.8 71.9 51.8 69.4
MB 25.6 25.5 76.5 53.2 73.4 12.4 35.4 67.4 49.7 65.7
chrF MC 29.5 28.5 77.6 53.8 74.4 17.3 48.9 72.8 53.1 70.4
MB 25.4 25.8 76.8 53.4 73.7 12.3 35.9 67.9 50.5 66.2
MC 28.1 26.5 84.1 55.4 78.7 15.8 46.7 77.3 54.1 73.0
MB 25.5 25.7 81.4 54.5 76.6 12.3 35.6 69.9 50.8 67.4
Bleurt MC OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
MB OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
RAMBR Bleu MC 29.8 27.7 77.3 53.7 74.2 17.2 46.8 72.1 52.0 69.6
MB 25.4 25.4 76.4 53.2 73.3 12.0 35.3 67.3 49.8 65.6
chrF MC 29.4 28.6 77.5 53.5 74.4 17.1 48.7 72.8 53.1 70.4
MB 25.1 25.7 76.6 53.0 73.6 12.2 35.8 67.8 50.5 66.2
Comet MC 28.7 26.9 83.2 55.4 78.0 16.0 46.9 76.6 53.9 72.6
MB 25.6 25.7 80.5 54.3 76.0 12.3 35.7 69.6 50.7 67.2
CBMBR Comet MC 27.6 26.0 84.0 55.2 78.5 15.7 46.5 77.2 53.7 72.8
MB 26.9 25.6 82.8 54.4 77.5 14.5 44.5 75.0 52.6 71.1
Pruning Bleu MC 29.8 27.8 77.3 53.8 74.2 17.2 46.8 72.0 51.8 69.4
MB 28.5 27.1 77.3 53.7 74.1 15.5 44.1 71.2 51.4 68.9
chrF MC 29.5 28.5 77.6 53.8 74.5 17.2 48.8 72.8 53.2 70.5
MB 28.8 27.4 77.6 53.9 74.4 15.9 45.2 71.7 52.2 69.4
Comet MC 28.1 26.5 84.1 55.5 78.6 15.9 46.8 77.3 54.1 72.9
MB 28.9 27.3 82.3 55.1 77.4 15.2 44.3 74.4 52.7 70.9
Bleurt MC 27.3 25.9 78.0 58.7 75.0 15.7 46.5 73.6 56.7 71.3
MB 28.3 26.9 77.7 55.7 74.6 15.1 44.2 72.0 53.7 69.7
PMBR Bleu MC 29.0 27.1 77.0 53.7 73.9 16.9 46.5 71.8 51.7 69.2
MB 28.9 27.0 76.7 53.2 73.7 16.5 45.8 71.4 51.4 68.8
chrF MC 29.0 27.9 77.3 53.7 74.2 16.0 47.6 72.2 52.5 69.8
MB 28.7 27.8 76.9 53.1 73.8 15.9 47.4 72.0 52.4 69.6
Comet MC 28.3 26.6 83.9 55.5 78.5 15.9 46.7 77.1 54.0 72.9
MB 27.0 26.1 82.5 54.9 77.4 14.1 41.5 73.5 52.3 70.0
Bleurt MC 27.6 26.2 78.0 58.6 74.9 15.5 46.4 73.5 56.6 71.2
MB 27.7 26.3 77.6 56.9 74.5 14.7 44.1 72.0 54.4 69.6
Oracle Bleu 45.5 41.9 80.9 61.4 74.2 31.9 54.9 74.4 56.3 69.0
chrF 45.6 42.8 81.0 61.3 74.1 28.2 58.2 75.5 57.8 70.0
Comet 35.4 33.4 87.6 61.8 79.3 20.5 50.8 81.2 58.9 73.9
Bleurt 36.3 34.2 81.9 68.6 75.6 21.3 50.9 76.6 63.8 71.6